# Route Sparse Autoencoder to Interpret Large Language Models

### v2

@@ 1 Introduction
Mechanistic interpretability of large language models (LLMs) seeks to understand and intervene in the internal process of information propagation and reasoning, to further improve trust and safety (Elhage et al., 2022b; Gurnee et al., 2023; Wang et al., 2023). Sparse autoencoders (SAEs) identify causally relevant and interpretable monosemantic features in LLMs, offering a promising solution for mechanistic interpretability (Bricken et al., 2023). Therefore, SAE and its variants (Huben et al., 2024; Rajamanoharan et al., 2024a; Gao et al., 2024; Rajamanoharan et al., 2024b) have been widely utilized in LLM interpretation tasks, such as feature discovery (Templeton et al., 2024; Gao et al., 2024) and circuit analysis (Marks et al., 2024).
@@
Typically, SAE is trained in an unsupervised manner. It first disentangles the intermediate activations from a single layer in the language model into a sparse, high-dimensional feature space, which is subsequently reconstructed by a decoder. This process reverses the effects of superposition (Elhage et al., 2022a) by extracting features that are sparse, linear, and decomposable. However, prior study (Yun et al., 2021) demonstrated that the activation strength of features in this sparse feature space exhibits distinct distribution patterns across layers 111Referred to as “Transformer factors” in (Yun et al., 2021)..
As illustrated in Figure 1, low-level features (e.g., “Crushed things” and “Europe”) representing word-level polysemy disambiguation peak in activation strength within shallow layers and gradually decline in deeper layers. Conversely, high-level features exhibit increasing activation strength as depth increases, corresponding to sentence-level or longer-range pattern formation, such as “Enumeration” and “One of the [number/quantifier]” 222Refer to (Yun et al., 2021) for more examples of low- and high-level features..
@@
This distribution disparity presents a significant challenge for previous SAEs (Huben et al., 2024; Rajamanoharan et al., 2024a; Gao et al., 2024; Rajamanoharan et al., 2024b), as they typically extract features from the hidden state of a single layer, failing to capture feature activating at other layers effectively (cf. Figure 2). Recently proposed Sparse Crosscoders (Lindsey et al., 2024)333Currently a conceptual framework without complete experimental validation. serve as an alternative to address this limitation, which separately encodes the hidden states of each layer into a high-dimensional feature space and aggregates the resulting representations for reconstruction (cf. Figure 2). This approach facilitates the joint learning of features across different layers. However, Crosscoder has two critical limitations: (1) Limited scalability: For an -layer model, Crosscoder employs separate encoders and decoders to process activations layer by layer, resulting in a parameter scale approximately times larger than traditional SAEs. This significantly increases computational overhead during both training and inference. (2) Uncontrollable interventions: Crosscoder’s joint learning mechanism limits the ability to adjust feature activations for specific LLM outputs in a controlled manner, reducing its flexibility for tasks that require precise feature-level intervention, e.g., feature steering (Templeton et al., 2024).
@@
To address these challenges, we propose a new framework — Route Sparse Autoencoder (RouteSAE). At the core is integrating a lightweight router with a shared SAE to dynamically extract multi-layer features in an efficient and flexible manner. A router is employed to compute normalized weights for activations from multiple layers. This dynamic weighting approach significantly reduces the number of parameters compared to a suite of layer-specific encoders and decoders, thereby addressing scalability concerns. Additionally, by unifying feature disentanglement and reconstruction within a shared SAE, RouteSAE facilitates fine-grained adjustments of specific feature activations, enabling more controlled interventions to influence the model’s output. This enhances flexibility and supports precise feature-level control, making the framework well-suited for tasks requiring robust and interpretable manipulation of model activations.
@@
We conducted comprehensive experiments on Llama-3.2-1B-Instruct (Dubey et al., 2024), evaluating downstream KL divergence, interpretable feature numbers, and interpretation score. The experimental results demonstrate that integrating a router into the SAE framework significantly improves the interpretability. At an equivalent sparsity level of 64, RouteSAE achieves a 22.5% increase in the number of interpretable features and a 22.3% improvement in interpretation scores. Additionally, we present a case study to highlight the practical advantages of the routing mechanism, illustrating how it improves feature interpretability by capturing features across multiple layers.
2 Related Work
In this section, we begin by reviewing prior work on sparse encoding, followed by a discussion of SAEs for interpreting LLMs. Finally, we briefly introduce works on cross-layer feature extraction in LLMs.
2.1 Sparse Encoding
Dictionary learning (Mairal et al., 2009) is a foundational machine learning approach that aims to learn an overcomplete set of basis components, enabling efficient data representation through sparse linear combinations. Autoencoders (Hinton & Salakhutdinov, 2006), in contrast, are designed to extract low-dimensional embeddings from high-dimensional data. By merging these two paradigms, sparse autoencoders have been developed, incorporating sparsity constraints such as regularization (Memisevic et al., 2015) to enforce sparsity in learned representations. Sparse autoencoders have found widespread application across various domains of machine learning, including computer vision (Wang et al., 2015) and natural language processing (Chang et al., 2018).
2.2 Sparse Autoencoder for LLMs
SAEs have emerged as effective tools for capturing monosemantic features (Elhage et al., 2022a), making them increasingly popular in LLM applications. Early work (Huben et al., 2024) introduced SAEs for extracting interpretable features from the internal activations of GPT-2 (Radford et al., 2019). To address systematic shrinkage in feature activations inherent in traditional SAEs (Huben et al., 2024; Bricken et al., 2023), Gated SAEs (Rajamanoharan et al., 2024a) were proposed, decoupling feature detection from magnitude estimation. TopK SAEs (Gao et al., 2024), inspired by k-sparse autoencoders (Makhzani & Frey, 2014), directly controlled sparsity to enhance reconstruction fidelity while preserving sparse representations. JumpReLU SAEs (Rajamanoharan et al., 2024b) advanced the trade-off between reconstruction quality and sparsity by replacing the conventional ReLU activation (Agarap, 2019) with the discontinuous JumpReLU function (Erichson et al., 2020). More recently, Switch SAEs (Mudide et al., 2024) introduced a mixture-of-experts mechanism, where inputs are routed to smaller, specialized SAEs, achieving better reconstruction performance within fixed computational constraints. However, these approaches capture the intermediate activations of language models from a single layer, neglecting features activated across multiple layers, which limits their overall applicability.
2.3 Features across Layers
Layer-wise differences in activation features within the transformer-based language model were first highlighted in (Yun et al., 2021), revealing that shallow layers capture low-level features while deeper layers focus on high-level patterns. Building on this, Gemma Scope (Lieberum et al., 2024) leveraged JumpReLU SAEs (Rajamanoharan et al., 2024b) to train separate models for each layer and sub-layer of the Gemma 2 models (Rivière et al., 2024). Similarly, Llama Scope (He et al., 2024) trained 256 SAEs per layer and sublayer of the Llama-3.1-8B-Base model (Dubey et al., 2024), extending layer-wise sparse modeling. Nevertheless, training a suite of SAEs is computationally expensive and often learns redundant features, posing significant scalability challenges for larger models. Moreover, determining the specific SAE relevant to a given input or characteristic can be nontrivial, complicating their practical application. Recently, Sparse Crosscoders (Lindsey et al., 2024) introduced a cross-layer SAE variant designed to investigate layer interactions and shared features (Templeton et al., 2024; Kissane et al., 2024). This framework facilitates circuit-level analysis (Elhage et al., 2021; Marks et al., 2024) by enabling feature tracking across layers, providing valuable insights into the evolution of model features and architectural differences. However, Crosscoder still relies on separate encoders and decoders for each layer, which limits its efficiency and hinders seamless integration with downstream tasks.
The challenges of scalability, feature localization, and applicability to downstream tasks identified in the works discussed in this section motivate the development of RouteSAE, a framework designed to overcome these limitations via a routing mechanism.
@@ 3 Methodology
In this section, we first provide a brief overview of SAEs, then introduce our proposed Route Sparse autoencoder (RouteSAE) in detail, highlighting its architecture, dynamic routing mechanism, and advantages in addressing the challenges of scalability and feature-level controllability.
@@ 3.1 Preliminary
SAE and Feature Decomposition. SAEs decompose language model activations — typically residual streams (He et al., 2016), , into a sparse linear combination of features , where represents the feature space dimension. The original activation is reconstructed using an encoder-decoder pair defined as follows:
@@
TopK SAE. Early SAEs (Huben et al., 2024; Bricken et al., 2023) leverage the ReLU activation function (Agarap, 2019) to generate sparse feature representations, coupled with an additional regularization term on latent representation to enforce sparsity. However, this approach is prone to feature shrinkage, where the constraint drives positive activations in toward zero, reducing the expressive capacity of the sparse feature space. To mitigate this issue, TopK SAE (Gao et al., 2024) replaces the ReLU activation function with a function, which directly controls the number of active latent dimensions by selecting the top largest values in . This is defined as:
By eliminating the need for an regularization term, TopK SAE achieves a more effective balance between sparsity and reconstruction quality, while enhancing the model’s ability to learn disentangled and interpretable monosementic features. In our RouteSAE framework, the shared SAE module is instantiated as a TopK SAE due to its superior performance in producing monosemantic features compared to other SAE variants.
3.2 Route Sparse Autoencoder
As shown in Figure 1, existing SAEs are typically trained on intermediate activations from a single layer, restricting their ability to simultaneously capture both low-level features from shallow layers and high-level features from deep layers. To overcome this limitation, we propose RouteSAE, which incorporates a lightweight router to dynamically integrate multi-layer residual stream activations from language models and disentangle them into a unified feature space.
Layer Weights. As illustrated in Figure 3, the router receives residual streams from multiple layers and determines which layer’s activation to route. Instead of concatenating these activations — an approach that could result in an excessively large input dimension — we adopt a simple yet effective aggregation strategy: sum pooling. Specifically, given activations from layer , we aggregate them using sum pooling to form the router’s input:
@@ 3.2 Route Sparse Autoencoder
The probability reflects the likelihood that the activation strength peaks at layer , dynamically assigned by the router based on the input representations.
@@
Shared SAE and Unified Feature Space. The routed intermediate activation ( , as defined in Equation 9) is processed by a shared SAE for reconstruction, which in this work is instantiated as a TopK SAE (Gao et al., 2024). Notably, RouteSAE is flexible and can be easily adapted to various SAE variants, including ReLU SAE (Huben et al., 2024; Bricken et al., 2023), Gated SAE (Rajamanoharan et al., 2024a) and JumpReLU SAE (Rajamanoharan et al., 2024b). By employing a shared SAE, RouteSAE establishes a unified feature space across activations from all routing layers. This ensures consistent and coherent feature representations, thereby enhancing the disentanglement of high-dimensional features and improving interpretability.
@@ 4 Experiments
In this section, we first outline the experimental setup, followed by the evaluation of RouteSAE. In this paper, we follow prior work (Gao et al., 2024; Rajamanoharan et al., 2024a; Huben et al., 2024; Templeton et al., 2024; He et al., 2024) and employ multiple evaluation metrics to assess the effectiveness of RouteSAE, including downstream KL-divergence, interpretable features, interpretation score, and reconstruction loss. Finally, we provide a detailed case study, demonstrating that RouteSAE effectively captures both low-level features from shallow layers and high-level features from deep layers.

### v3

@@ Abstract
Route Sparse Autoencoder to Interpret Large Language Models
Wei Shi1††thanks: Equal Contribution, Sihang Li111footnotemark: 1, Tao Liang2, Mingyang Wan2, Guojun Ma2††thanks: Corresponding, Xiang Wang122footnotemark: 2, Xiangnan He1, 1University of Science and Technology of China, 2Douyin Co., Ltd., swei2001@mail.ustc.edu.cn, taoliangdpg@126.com, {wanmingyang, maguojun}@bytedance.com, {sihang0520, xiangwang1223, xiangnanhe}@gmail.com
@@ 1 Introduction
Mechanistic interpretability of large language models (LLMs) seeks to understand and intervene in the internal process of information propagation and reasoning, to further improve trust and safety Elhage et al. (2022b); Gurnee et al. (2023); Wang et al. (2023). Sparse autoencoders (SAEs) identify causally relevant and interpretable monosemantic features in LLMs, offering a promising solution for mechanistic interpretability Bricken et al. (2023). Therefore, SAE and its variants Huben et al. (2024); Rajamanoharan et al. (2024a); Gao et al. (2024); Rajamanoharan et al. (2024b) have been widely utilized in LLM interpretation tasks, such as feature discovery Templeton et al. (2024); Gao et al. (2024) and circuit analysis Marks et al. (2024).
@@
Typically, SAE is trained in an unsupervised manner. It first disentangles the intermediate activations from a single layer in the language model into a sparse, high-dimensional feature space, which is subsequently reconstructed by a decoder. This process reverses the effects of superposition Elhage et al. (2022a) by extracting features that are sparse, linear, and decomposable.
However, the activation strength of features in this feature space exhibits distinct distribution patterns across layers111Referred to as “Transformer factors” in Yun et al. (2021). Yun et al. (2021). As shown in Figure 1, low-level features, which are associated with disambiguating word-level polysemy, tend to exhibit peak activation in the early layers and decline steadily in deeper layers. High-level features, which capture sentence-level or long-range structure, show increasing activation with depth.222Refer to Yun et al. (2021) for more examples of low- and high-level features.
@@
This distribution disparity presents a significant challenge for previous SAEs Huben et al. (2024); Rajamanoharan et al. (2024a); Gao et al. (2024); Rajamanoharan et al. (2024b), as they typically extract features from the hidden state of a single layer, failing to capture feature activating at other layers effectively (cf. Figure 2). Recently proposed Sparse Crosscoders333Currently a conceptual framework without complete experimental validation. Lindsey et al. (2024) serve as an alternative to address this limitation, which separately encodes the hidden states of each layer into a high-dimensional feature space and aggregates the resulting representations for reconstruction (cf. Figure 2). This approach facilitates the joint learning of features across different layers. However, Crosscoder has two critical limitations: (1) Limited scalability: For an -layer model, Crosscoder employs separate encoders and decoders to process activations layer by layer, resulting in a parameter scale approximately times larger than traditional SAEs. This significantly increases computational overhead during both training and inference. (2) Uncontrollable interventions: Crosscoder’s joint learning mechanism projects hidden states into a high-dimensional space and then aggregates them, making it impractical to precisely identify and adjust the feature activations at specific layers. This limits its flexibility for tasks requiring controlled, feature-level interventions, e.g., feature steering Templeton et al. (2024).
@@
To address these challenges, we propose Route Sparse Autoencoder (RouteSAE). At the core is integrating a lightweight router with a shared SAE to dynamically extract multi-layer features in an efficient and flexible manner. A router is employed to compute normalized weights for activations from multiple layers. This dynamic weighting approach significantly reduces the number of parameters compared to a suite of layer-specific encoders and decoders, thereby addressing scalability concerns. Additionally, by unifying feature disentanglement and reconstruction within a shared SAE, RouteSAE facilitates fine-grained adjustments of specific feature activations, enabling more controlled interventions to influence the model’s output. This enhances flexibility and supports precise feature-level control, making the framework well-suited for tasks requiring robust and interpretable manipulation of model activations.
@@
We conduct comprehensive experiments on Llama-3.2-1B-Instruct Dubey et al. (2024), evaluating downstream KL divergence, interpretable feature numbers, and interpretation score. The experimental results demonstrate that RouteSAE significantly improves the interpretability. At an equivalent sparsity level of 64, it achieves a 22.5% increase in the number of interpretable features and a 22.3% improvement in interpretation scores.
Our contributions are summarized as follows:
We propose RouteSAE, a novel sparse autoencoder framework that integrates multi-layer activations through a routing mechanism.
RouteSAE achieves higher computational efficiency than Crosscoder by using a shared SAE structure with minimal additional parameters.
Extensive experiments confirm that RouteSAE enhances model interpretability, highlighting the effectiveness of the proposed routing mechanism.
@@ 2 Methodology
In this section, we first briefly review SAEs, then introduce our proposed Route Sparse Autoencoder (RouteSAE) in detail.
@@ 2.1 Preliminary
SAE and Feature Decomposition. SAEs decompose language model activations — typically residual streams He et al. (2016), , into a sparse linear combination of features , where represents the feature space dimension. The original activation is reconstructed using an encoder-decoder pair defined as follows:
@@
TopK SAE. Early SAEs Huben et al. (2024); Bricken et al. (2023) leverage the ReLU activation function Agarap (2019) to generate sparse feature representations, coupled with an additional regularization term on latent representation to enforce sparsity. However, this approach is prone to feature shrinkage, where the constraint drives positive activations in toward zero, reducing the expressive capacity of the sparse feature space. To mitigate this issue, TopK SAE Gao et al. (2024) replaces the ReLU activation function with a function, which directly controls the number of active latent dimensions by selecting the top largest values in . This is defined as:
By eliminating the need for an regularization term, TopK SAE achieves a more effective balance between sparsity and reconstruction quality, while enhancing the model’s ability to learn disentangled and interpretable monosementic features. In our RouteSAE framework, the shared SAE module is instantiated as a TopK SAE due to its superior performance in producing monosemantic features.
2.2 Route Sparse Autoencoder (RouteSAE)
As shown in Figure 2, existing SAEs are typically trained on intermediate activations from a single layer, restricting their ability to simultaneously capture both low-level features from shallow layers and high-level features from deep layers. To overcome this limitation, RouteSAE incorporates a lightweight router to dynamically integrate multi-layer residual streams from language models and disentangle them into a unified feature space.
Layer Weights. As illustrated in Figure 3, the router receives residual streams from multiple layers and determines which layer’s activation to route. Instead of concatenating these activations, which could result in an excessively large input dimension, we adopt a simple yet effective aggregation strategy: sum pooling. Specifically, given activations from layer , we aggregate them using sum pooling to form the router’s input:
@@ 2.2 Route Sparse Autoencoder (RouteSAE)
reflects the likelihood that the activation strength peaks at layer , dynamically assigned by the router based on the input representations.
@@
Shared SAE and Unified Feature Space. The routed intermediate activation ( , as defined in Equation 9) is processed by a shared SAE for reconstruction, which in this work is instantiated as a TopK SAE Gao et al. (2024). Notably, RouteSAE is flexible and can be easily adapted to various SAE variants. By employing a shared SAE, RouteSAE establishes a unified feature space across activations from all routing layers. This ensures consistent feature representations, thereby enhancing the disentanglement of high-dimensional features and improving interpretability.
@@ 3 Experiments
We first outline the experimental setup, followed by the evaluation of RouteSAE. In this paper, we follow prior work Gao et al. (2024); Rajamanoharan et al. (2024a); Huben et al. (2024); Templeton et al. (2024); He et al. (2024) and employ multiple evaluation metrics to assess the effectiveness of RouteSAE, including downstream KL-divergence, interpretable features, interpretation score, and reconstruction loss. Finally, we provide a detailed case study, demonstrating that RouteSAE not only effectively captures both low-level features from shallow layers and high-level features from deep layers, but also enables targeted manipulation of these features to control the model’s output.
@@ 3.1 Setup
Inputs. We train all SAEs on the residual streams of the Llama-3.2-1B-Instruct. For baseline SAEs, we follow the standard approach Gao et al. (2024) of selecting the layer located approximately at of the model depth (i.e., Layer 11). Prior work Lad et al. (2024) has shown that the early layers of LLMs primarily handle detokenization, whereas later layers specialize in next-token prediction. Based on this insight, we select residual streams from the middle layers of the model as input for both RouteSAE and Crosscoder Lindsey et al. (2024). In particular, we focus on layers spanning to of the model depth, as detailed in Table 1.
The training data is sourced from OpenWebText2 Gao et al. (2020), comprising approximately 100 million randomly sampled tokens for training, with an additional 10 million tokens reserved for evaluation. All experiments are conducted using a context length of 512 tokens. To ensure stable training, we normalize the language model activations following Gao et al. (2024).
Hyperparameters. For all SAEs, we use the Adam optimizer Kingma and Ba (2015) with standard settings: and . The learning rate is set to , following a three-phase schedule. (1) Linear warmup. The learning rate increases linearly from 0 to the target rate over the first 5% of training steps. (2) Stable phase. The learning rate remains constant for 75% of the training steps. (3) Linear Decay. The learning rate gradually decreases to zero over the final 20% of training steps to ensure smooth convergence. To improve training stability, we apply unit norm regularization Gao et al. (2024) to the columns of the SAE decoder every 10 steps, ensuring that the decoder columns maintain unit length.
Baselines. We benchmark RouteSAE against leading baselines, including ReLU SAE Huben et al. (2024), Gated SAE Rajamanoharan et al. (2024a), TopK SAE, and Crosscoder Lindsey et al. (2024). Moreover, we compare with a random setting, where the router is replaced by a uniform distribution that assigns equal routing weights to each layer. It is important to note that Crosscoder remains a conceptual framework and lacks complete experimental validation. As there is no official codebase or hyperparameter guidance available, we implement it following the description in Lindsey et al. (2024). We acknowledge that our results may not fully reflect its actual performance.
3.2 Downstream KL Divergence
To assess whether the extracted features are relevant for language modeling, we replace the residual streams with the reconstructed representation during the forward pass of the language model and evaluate the reconstruction quality using Kullback-Leibler (KL) divergence. It quantifies the discrepancy between the original and reconstructed distributions, with lower value indicating that the extracted features are highly relevant for language modeling. Note that RouteSAE replaces the activation at the layer with the highest routing weight.
As shown in Figure 4, the sparsity-KL divergence frontiers for ReLU and Gated SAE are nearly identical, yet both exhibit a significant gap compared to TopK SAE. Due to suboptimal reconstruction quality, the KL divergence for ReLU and Gated SAE drops substantially as increases, falling from around 400 to 350. In contrast, the KL divergence for both TopK and RouteSAE remains consistently below 150, with only minimal decreases as increases. This indicates that both methods are able to effectively reconstruct the original input even at high sparsity levels. The random routing baseline yields higher KL divergence than both TopK and RouteSAE, further highlighting the advantage of learned routing.
Notably, RouteSAE achieves the best performance among all methods, maintaining a lower KL divergence at comparable sparsity levels. It outperforms even TopK SAE, indicating that feature substitution during inference is most effective when performed at the layer where the target feature is most active, rather than at a predetermined fixed layer. We exclude Crosscoder from this comparison, as it produces multiple reconstructed representations , making its application to this setting nontrivial and not directly comparable.
3.3 Interpretable Features
Previous works Huben et al. (2024); He et al. (2024) interpret features by preserving the context with the highest feature activation value. However, we argue that it has two limitations: (1) Retaining only the highest activation context for each feature leads to a large number of undiscernible features; (2) Each feature is associated with only a single context, reducing the reliability of the interpretation.
To address these limitations, we introduce a new approach for preserving feature contexts using an activation threshold. For a given sequence context, only features with activation values exceeding the threshold are retained. As shown in Figure 5(a), increasing the threshold reduces the number of retained features. In contrast, Figure 5(b) demonstrates that a higher threshold leads to improved interpretation scores. Consequently, the threshold governs a trade-off between the quantity of interpretable features and their interpretability quality. In this section, we set the threshold to 15, which achieves a balance between maintaining sufficient feature quantity and enhancing interpretability. Notably, a single sequence may be associated with multiple contexts.
As illustrated in Figure 6, at a threshold of 15, both ReLU and Gated SAE extract over 1,000 interpretable features, performing similarly. In contrast, TopK SAE significantly outperforms both, extracting more than 3,000 features. RouteSAE surpasses all other methods, extracting over 4,000 features at the same threshold. Notably, RouteSAE exhibits a more gradual decline in the number of extracted features as increases, while TopK SAE exhibits a more pronounced reduction. The random routing baseline sometimes extracts even more features than RouteSAE, but its feature count decreases much more rapidly as increases. These results suggest that learning based solely on single-layer activation values limits the ability of SAEs to extract interpretable features. In comparison, Crosscoder extracts substantially fewer features, retaining approximately 200. Since Crosscoder aggregates and projects activations across multiple layers, we hypothesize that the optimal threshold for balancing feature quantity and interpretability lies in a lower range for Crosscoder. Therefore, comparing it against the same activation threshold may not reflect its actual ability to extract high-quality features. We plan to investigate this in future work.
3.4 Interpretation Score
Despite the feature screening in Section 3.3, the number of retained features remains in the thousands, making manual interpretation and evaluation challenging. To further assess feature interpretability, we follow prior work Huben et al. (2024); Templeton et al. (2024); He et al. (2024) and leverage GPT-4o Hurst et al. (2024) to analyze the features extracted by SAEs, assigning an interpretability score alongside feature descriptions. Unlike previous approaches, we provide GPT-4o with multiple token categories per feature along with their contextual usage. Given resource constraints, we randomly select a subset of 100 retained features per SAE for interpretation. As detailed in Appendix D, for each feature, we construct a structured prompt comprising a prefix prompt, the activated token, and its surrounding context, which is then given to GPT-4o. GPT-4o outputs three standardized components: (1) Feature categorization, labeling each feature as low-level, high-level, or undiscernible; (2) Interpretability score, rated on a scale of 1 to 5; and (3) Explanation, providing a brief justification for the assigned category and score.
To evaluate the consistency between GPT-4o and human annotators in both feature categorization and interpretability scoring, we randomly sample 100 features from RouteSAE. For each feature, we provide its activation contexts and a scoring prompt to both GPT-4o and human annotators. As illustrated in Figure 7, (a) shows the percentage of features assigned to each interpretability category (“Low,” “High,” or “Undiscernible”) by both humans and GPT-4o. The two distributions are nearly identical, reflecting strong categorical agreement between human and GPT-4o annotations. (b) depicts the distribution of absolute differences in interpretability scores, showing that most features exhibit minimal discrepancy between human and GPT-4o (). This indicates a high degree of alignment in interpretability assessment.
Figure 8 shows that both ReLU and Gated SAE exhibit low and relatively stable interpretation scores, consistently falling below those of the other methods. TopK SAE shows a noticeable decline in interpretation scores as increases, with scores dropping from over 4.0 at sparsity 48 to around 3.7 at sparsity 72. In contrast, Crosscoder, despite not being sensitive to changes in sparsity, maintains consistent scores, hovering around 3.9 across all sparsity levels. The random routing baseline achieves higher interpretation scores than ReLU, Gated, TopK, and Crosscoder, but remains consistently lower than RouteSAE. In comparison, RouteSAE achieves the highest interpretation scores, maintaining values above 4.4 at all sparsity levels. It remains largely unaffected by changes in sparsity, demonstrating its robust ability to preserve high interpretability, regardless of the sparsity setting. These results indicate that dynamically leveraging multi-layer activations, as done in RouteSAE and even to some extent by the random router, not only allows for extraction of more features but also leads to higher feature interpretability.
3.5 Routing Weights
In fact, reconstructing the activations of a language model using SAE becomes increasingly difficult as the number of layers grows. This is likely due to the increasing abstraction and entanglement of features in deeper layers, which imposes additional challenges on the autoencoder’s capacity to isolate and reconstruct meaningful components.
To analyze how RouteSAE allocates routing weights across layers during training and inference, we track the layer-wise routing weights throughout both phases. As shown in Figure 9, RouteSAE produces a distinct weight profile across layers, exhibiting a U-shaped distribution rather than concentrating weights on a small subset of shallow layers. This pattern suggests a balanced allocation of representational capacity, where both shallow and deep layers contribute meaningfully. These results are consistent with the observations reported in Yun et al. (2021), which indicate that lower-level features are primarily activated in the earlier layers, whereas higher-level features become more prominent in the deeper layers.
3.6 Case Study
Interpretable Features. As shown in Figure 3, RouteSAE effectively captures both low-level and high-level features from shallow and deep layers, respectively. Specifically, RouteSAE identifies low-level features such as “units of weight” and “Olympics” from shallow layers. The “units of weight” feature activates on tokens related to weight units, including terms like “pound” and “kilograms”. The “Olympics” feature captures variations of the term “Olympic”, such as “Olympics” and “Olympian”. These two features exemplify word-level polysemy disambiguation, peaking at shallow layers. At deeper layers, RouteSAE extracts high-level features, including the patterns “more [X] than [Y]” and “do everything [possible/in my power]”. The first feature identifies tokens that appear in comparative structures, particularly those following the pattern “more [X] than [Y].” The second feature highlights tokens in phrases expressing a commitment to maximal effort or capability, such as “do my best”, and “do all he could”. These two features reflect sentence-level or long-range pattern formation, peaking at deeper layers. These observations demonstrate that RouteSAE successfully integrates features from multiple layers of activations into a unified feature space. For more interpretable features, refer to Appendix E.
RouteSAE Feature Steering
Figure 10 illustrates how RouteSAE enables controlled model steering by directly manipulating internal features from the SAE decoder. This is achieved by replacing the activation with the reconstructed representation . In each example, the original response is generated without intervention, reflecting the model’s default behavior. In contrast, the clamped response is obtained by increasing the activation of a specific target feature to 20. In the upper example, the clamped feature is a low-level one related to the “Olympics” concept; after intervention, the model’s response becomes focused on Olympic-related content, regardless of the input question. In the lower example, the manipulated feature is a high-level one representing the intent to “do everything possible”; as a result, the model adopts a proactive, determined stance, as evidenced by responses such as “I can do this.” This illustrates that, RouteSAE enables more controllable and targeted interventions on model behavior through direct feature activation manipulation.
4 Conclusion
Limitations
Ethical Considerations
Appendix A Related Work
In this section, we begin by reviewing prior work on sparse encoding, followed by a discussion of SAEs for interpreting LLMs. Finally, we briefly introduce cross-layer feature extraction in LLMs.
A.1 Sparse Encoding
Dictionary learning Mairal et al. (2009) is a foundational machine learning approach that aims to learn an overcomplete set of basis components, enabling efficient data representation through sparse linear combinations. Autoencoders Hinton and Salakhutdinov (2006), in contrast, are designed to extract low-dimensional embeddings from high-dimensional data. By merging these two paradigms, sparse autoencoders have been developed, incorporating sparsity constraints such as regularization Memisevic et al. (2015) to enforce sparsity in learned representations. Sparse autoencoders have found widespread application across various domains of machine learning, including computer vision Wang et al. (2015) and natural language processing Chang et al. (2018).
A.2 Sparse Autoencoder for LLMs
SAEs have emerged as effective tools for capturing monosemantic features Elhage et al. (2022a), making them increasingly popular in LLM applications. Early work Huben et al. (2024) introduced SAEs for extracting interpretable features from the internal activations of GPT-2 Radford et al. (2019). To address systematic shrinkage in feature activations inherent in traditional SAEs Huben et al. (2024); Bricken et al. (2023), Gated SAEs Rajamanoharan et al. (2024a) were proposed, decoupling feature detection from magnitude estimation. TopK SAEs Gao et al. (2024), inspired by k-sparse autoencoders Makhzani and Frey (2014), directly controlled sparsity to enhance reconstruction fidelity while preserving sparse representations. JumpReLU SAEs Rajamanoharan et al. (2024b) advanced the trade-off between reconstruction quality and sparsity by replacing the conventional ReLU activation Agarap (2019) with the discontinuous JumpReLU function Erichson et al. (2020). More recently, Switch SAEs Mudide et al. (2024) introduced a mixture-of-experts mechanism, where inputs are routed to smaller, specialized SAEs, achieving better reconstruction performance within fixed computational constraints. However, these approaches capture the intermediate activations of language models from a single layer, neglecting features activated across multiple layers, which limits their overall applicability.
A.3 Features across Layers
Layer-wise differences in activation features within the transformer-based language model were first highlighted in Yun et al. (2021), revealing that shallow layers capture low-level features while deeper layers focus on high-level patterns. Building on this, Gemma Scope Lieberum et al. (2024) leveraged JumpReLU SAEs Rajamanoharan et al. (2024b) to train separate models for each layer and sub-layer of the Gemma 2 models Rivière et al. (2024). Similarly, Llama Scope He et al. (2024) trained 256 SAEs per layer and sublayer of the Llama-3.1-8B-Base model Dubey et al. (2024), extending layer-wise sparse modeling. Nevertheless, training a suite of SAEs is computationally expensive and often learns redundant features, posing significant scalability challenges for larger models. Moreover, determining the specific SAE relevant to a given input or characteristic can be nontrivial, complicating their practical application. Recently, Sparse Crosscoders Lindsey et al. (2024) introduced a cross-layer SAE variant designed to investigate layer interactions and shared features Templeton et al. (2024); Kissane et al. (2024). This framework facilitates circuit-level analysis Elhage et al. (2021); Marks et al. (2024) by enabling feature tracking across layers, providing valuable insights into the evolution of model features and architectural differences. However, Crosscoder still relies on separate encoders and decoders for each layer, which limits its efficiency and hinders seamless integration with downstream tasks.
The challenges of scalability, feature localization, and applicability to downstream tasks motivate the development of RouteSAE.
Appendix B Comparison of Routing Mechanisms.
Discussion. Hard routing enforces sparsity by selecting the activation from a single layer, typically the one with the strongest response for a given input. This mechanism simplifies the routing task, as the router only needs to identify the layer with the highest activation. In contrast, soft routing aggregates activations from all layers, weighted by their estimated importance scores. This introduces a significantly more challenging requirement: the router must accurately estimate the relative contribution of each layer. Inaccurate estimations may result in disproportionately high weights assigned to less relevant layers, which can lead to the accumulation of noisy or irrelevant activations. This, in turn, may interfere with the disentanglement of monosemantic features in subsequent stages. While soft routing has the potential to capture cross-layer features—i.e., features that are distributed across multiple layers—our experiments thus far have not demonstrated clear benefits in this setting. We plan to investigate this direction further in future work.